Service fully restored at 19:19 after a hardware change at Faraday. Incident closed following a successful monitoring period.
Update · UTC
A replacement amplifier card has been delivered and a second is on the way. An engineer is on route to replace the failed amplifier/card. ETA to follow. The estimated time of restoration is not yet known. Next update will be issued upon service restoration unless there should be a significant development.
Identified · UTC
Customer Impact : Loss of service to 70 Optical Connect circuits.
Current Situation : The cold restart performed earlier, failed to restore service and a replacement has been ordered. ETA 16:30. An engineer will be tasked in due course. The estimated time of restoration is not yet known. Next update will be issued upon service restoration unless there should be a significant development.